OpenStudy (anonymous):

Mary and her brother John collect foreign coins. Mary has three times the number of coins that John has. Together they have 180 foreign coins. Find how many coins Mary has.

OpenStudy (anonymous):

M = 3J M + J = 180 3J + J = 180 4J = 180 J = 45 M = 3(45) M = 135 coins